# Quantum fluctuation-induced first-order breaking of time-reversal symmetry in unconventional superconductors

### v1

@@ Abstract
Figure 4 shows the phase stiffness without and with the phase fluctuations. The stiffness tensor for the square lattice satisfies $D_{\alpha\beta}=D_{xx}\delta_{\alpha\beta}$. The mean-field stiffness $D_{xx}^{(0)}$ is a continuous function of the chemical potential and temperature, whereas the phase-fluctuating stiffness $D_{xx}$ is discontinuous, reflecting again the first-order nature of the quantum phase transition. Moreover, the phase fluctuations suppress the phase stiffness even at zero temperature around the $s+id$ dome, i.e., near the quantum critical point. The maximum suppression at zero temperature within the range of chemical potential in Fig. 4 is $15\%$. We find that there is almost no stiffness suppression for chemical potentials far away from the quantum critical point. Given that our system is clean, this result recognizes the important role of the $\mathcal{T}$-breaking quantum critical point in suppressing the phase stiffness in the ground state.
@@
Figure 4: Phase stiffness without [left panel, denoted with superscript $(0)$] and with (right panel) the phase fluctuations as functions of the chemical potential and temperature around the $s+id$ dome.
@@
The formalism we presented is essentially a self-consistent Gaussian approximation for the phase fluctuations that couple to the long-range charge fluctuations on the level of the random phase approximation. Because it is a self-consistent approach, this formalism takes into account some higher-order effects of the phase fluctuations beyond the bare Gaussian approximation. It should be these higher-order effects that modify the order of the $\mathcal{T}$-breaking quantum phase transition.
Increasing the ratio of the cutoff wavevector over the gap amplitude $q_{c}/|\Delta_{x}|$ within a limit leads to a more profound modification of the phase diagram but keeps the qualitative features presented in this work. However, if $q_{c}/|\Delta_{x}|$ is set beyond the limit, e.g., $q_{c}=4|\Delta_{x}|/v_{F}$, we find that the saddle point solution will be unstable at high temperatures well above the superconducting transition temperature [29].
@@
Finally, the formalism presented in this work can be extended straightforwardly for other systems such as twisted cuprate Josephson junctions promising to host high-temperature topological $d_{x^{2}-y^{2}}+id_{xy}$ superconductivity [16, 19], which will involve multiple bands and will be a future work. Although we have not directly calculated for this twist-induced $d_{x^{2}-y^{2}}+id_{xy}$ phase, our current findings may still provide insights for interpreting the recent experiment [19] on it. The experiment showed an interesting feature: the reversible Josephson diode effect implying breaking $\mathcal{T}$ has a complex dependence on the twist angle and sharply vanishes for some twist angles within $(40^{\circ},44^{\circ})$ (Fig. 4F, G in Ref. [19]). This suggests that the $d_{x^{2}-y^{2}}+id_{xy}$ phase does not exist for some twist angles where the mean-field theory predicts to support it [16] and that the associated quantum phase transition is first order. It is, however, consistent with our findings of the splitting of the $\mathcal{T}$-broken superconducting phase region and the modification of phase-transition order near the quantum critical point, which result from the quantum phase fluctuations. Therefore, it is possible that our findings point out some general effects of the phase fluctuations on the spontaneous $\mathcal{T}$ breaking in superconductors, which could shape our understanding of such important quantum phase transitions and have implications for the $\mathcal{T}$-broken topological superconductivity.
@@ Appendix A End Matter
$S=\int_{0}^{1/T}d\tau\biggl\{-\sum_{\bm{rr}^{\prime}}\psi_{\bm{r}}^{\dagger}(\tau)\bigl[\mathcal{G}^{-1}(\bm{r}-\bm{r}^{\prime},\tau)-\Sigma(\bm{r},\bm{r}^{\prime},\tau)\bigr]\psi_{\bm{r}^{\prime}}(\tau)-\sum_{\bm{q}}\frac{|\phi_{\bm{q}}(\tau)|^{2}}{2NV_{\bm{q}}}\biggr\}+\sum_{\langle\bm{rr}^{\prime}\rangle}\frac{2|\Delta_{\bm{r}-\bm{r}^{\prime}}|^{2}}{TJ},$ (7)
@@
$\Sigma(\bm{r},\bm{r}^{\prime},\tau)$ $=\Bigl\{\bigl[\mathrm{i}\partial_{\tau}\theta_{\bm{r}}(\tau)+\phi_{\bm{r}}(\tau)\bigr]\delta_{\bm{rr}^{\prime}}-t_{\bm{r}-\bm{r}^{\prime}}\bigl[\mathrm{e}^{-\mathrm{i}(\theta_{\bm{r}}(\tau)-\theta_{\bm{r}^{\prime}}(\tau))\tau_{3}}-1\bigr]\Bigr\}\tau_{3},$ (9)
@@
$+\frac{2}{NTJ}\sum_{\bm{k}\bm{q}}\bigl[\cos(k_{x}-q_{x})+\cos(k_{y}-q_{y})\bigr]\Delta_{\bm{k}}\Delta_{\bm{q}}^{*}-\sum_{\bm{q}n}\frac{|\phi_{\bm{q}n}|^{2}}{2NTV_{\bm{q}}},$ (10)
@@
$S_{\mathrm{eff}}=-\mathrm{Tr}\bigl[\ln\bigl((T\mathcal{G})^{-1}\bigr)\bigr]+\sum_{\gamma=1}^{\infty}\frac{1}{\gamma}\mathrm{Tr}\bigl[(\mathcal{G}\Sigma)^{\gamma}\bigr]+\frac{2N}{TJ}\bigl(|\Delta_{x}|^{2}+|\Delta_{y}|^{2}\bigr)-\sum_{\bm{q}n}\frac{|\phi_{\bm{q}n}|^{2}}{2NTV_{\bm{q}}},$ (11)
@@
$\Sigma(\bm{k},\bm{q},\mathrm{i}\nu_{n})\approx\frac{1}{\sqrt{N}}\bigl[\nu_{n}\tau_{3}+\mathrm{i}(\xi_{\bm{k}}-\xi_{\bm{q}})\bigr]\theta_{\bm{k}-\bm{q},n}-\frac{1}{2N}\sum_{\bm{k}^{\prime}m}(\xi_{\bm{k}}+\xi_{\bm{q}}-2\xi_{\bm{k}-\bm{k}^{\prime}})\tau_{3}\theta_{\bm{k}^{\prime}m}\theta_{\bm{k}-\bm{q}-\bm{k}^{\prime},n-m}+\frac{1}{N}\phi_{\bm{k}-\bm{q},n}\tau_{3}.$ (12)
@@
$\mathrm{Tr}\bigl[(\mathcal{G}\Sigma)^{2}\bigr]\approx{}$ $\frac{1}{N}\sum_{\bm{k}\bm{q}mn}\mathrm{Tr}\Bigl[\mathcal{G}(\bm{k},\mathrm{i}\omega_{n})\bigl[\bigl(\nu_{m}\theta_{\bm{q}m}+N^{-1/2}\phi_{\bm{q}m}\bigr)\tau_{3}+\mathrm{i}(\xi_{\bm{k}}-\xi_{\bm{k}-\bm{q}})\theta_{\bm{q}m}\bigr]$
$\times\mathcal{G}(\bm{k}-\bm{q},\mathrm{i}\omega_{n-m})\bigl[\bigl(-\nu_{m}\theta_{-\bm{q},-m}+N^{-1/2}\phi_{-\bm{q},-m}\bigr)\tau_{3}-\mathrm{i}(\xi_{\bm{k}}-\xi_{\bm{k}-\bm{q}})\theta_{-\bm{q},-m}\bigr]\Bigr].$ (14)
Since $\theta_{\bm{r}}(\tau)$ is varying slowly in space and time, $\theta_{\bm{q}n}$ would quickly vanish as $\bm{q}$ and $n$ depart from zero. We can thus further expand Eqs. (13, 14) in powers of $\bm{q}$ and $\nu_{n}$ up to their quadratic order because they are controlled by $\theta_{\bm{q}n}$,
$\mathrm{Tr}[\mathcal{G}\Sigma]$ $\approx\frac{1}{2T}\sum_{\bm{q}n}\sum_{\alpha,\beta=x}^{z}D^{\prime}_{\alpha\beta}q_{\alpha}q_{\beta}|\theta_{\bm{q}n}|^{2},$ (15)
$\mathrm{Tr}\bigl[(\mathcal{G}\Sigma)^{2}\bigr]$ $\approx\frac{1}{T}\sum_{\bm{q}n}\biggl[-\chi\biggl(\nu_{n}\theta_{\bm{q}n}+\frac{\phi_{\bm{q}n}}{\sqrt{N}}\biggr)\biggl(-\nu_{n}\theta_{\bm{q}n}^{*}+\frac{\phi_{\bm{q}n}^{*}}{\sqrt{N}}\biggr)+\sum_{\alpha,\beta=x}^{z}D^{\prime\prime}_{\alpha\beta}q_{\alpha}q_{\beta}|\theta_{\bm{q}n}|^{2}\biggr],$ (16)
@@
Here $n_{F}(\cdot)$ is the Fermi distribution function and $n_{F}^{\prime}(\cdot)$ represents its derivative. We can integrate Eq. (17) by parts and use the periodicity of $\xi_{\bm{k}}$ to get the phase stiffness
@@
Finally, plugging Eqs. (15, 16) into Eq. (11) and integrating out the fields $\phi_{\bm{q}n}$ and $\theta_{\bm{q}n}$, we obtain the free energy as in Eq. (3).

### v2

@@ Abstract
Figure 4 shows the phase stiffness without and with the phase fluctuations. The stiffness tensor for the square lattice satisfies $D_{\alpha\beta}=D_{xx}\delta_{\alpha\beta}$. The mean-field stiffness $D_{xx}^{(0)}$ is a continuous function of the chemical potential and temperature, whereas the phase-fluctuating stiffness $D_{xx}$ is discontinuous, reflecting again the first-order nature of the quantum phase transition. Moreover, the phase fluctuations suppress the phase stiffness even at zero temperature around the $s+\mathrm{i}d$ dome, i.e., near the quantum critical point. We find that there is almost no stiffness suppression for chemical potentials far away from the quantum critical point. Given that our system is clean, this result recognizes the important role of the $\mathcal{T}$-breaking quantum critical point in suppressing the phase stiffness in the ground state.
@@
Figure 4: Phase stiffness without [left panel, denoted with superscript $(0)$] and with (right panel) the phase fluctuations as functions of the chemical potential and temperature around the $s+\mathrm{i}d$ dome.
@@
The formalism we presented is essentially a self-consistent (because we minimized the total free energy that includes the phase-fluctuation correction) Gaussian approximation for the phase fluctuations that couple to the long-range charge fluctuations at the level of the random phase approximation. Because it is a self-consistent approach, this formalism takes into account some higher-order effects of the phase fluctuations beyond the bare Gaussian approximation. It should be these higher-order effects that modify the order of the $\mathcal{T}$-breaking quantum phase transition.
Increasing the cutoff wavevector within a limit leads to a more profound modification of the phase diagram but retains the qualitative features presented in this work. However, if $q_{c}/|\Delta_{x}|$ is set beyond the limit, e.g., $q_{c}=4|\Delta_{x}|/v_{F}$, we find that the saddle point solution will be unstable at high temperatures well above the superconducting transition temperature [29].
@@
Finally, the formalism presented in this work can be extended straightforwardly for other systems such as twisted cuprate Josephson junctions promising to host high-temperature topological $d_{x^{2}-y^{2}}+\mathrm{i}d_{xy}$ superconductivity [16, 19]. In this case, the only extension is to include multiple orbitals because of the enlarged unit cell of the bilayer system. Due to the similar mathematical structures of the formulations, although we have not directly calculated for this twist-induced $d_{x^{2}-y^{2}}+\mathrm{i}d_{xy}$ phase, our current findings may still provide insights for interpreting the recent experiment [19] on it. The experiment showed an intricate feature: the reversible Josephson diode effect, which implies breaking $\mathcal{T}$, sharply vanishes for some twist angles within $(40^{\circ},44^{\circ})$ (Fig. 4F, G in Ref. [19]). This suggests that the $d_{x^{2}-y^{2}}+\mathrm{i}d_{xy}$ phase does not exist for some twist angles where the mean-field theory predicts to support it [16] and that the associated quantum phase transition is of first order. It is, however, consistent with our findings of the phase fluctuation-induced splitting of the $\mathcal{T}$-broken phase region and modification of the phase-transition order near the quantum critical point. It will be a promising direction to use our formalism to directly study chiral superconductivity in twisted cuprate Josephson junctions and to make quantitative comparisons with experimental data.
@@ Appendix A End Matter
$S=\int_{0}^{1/T}\mathrm{d}\tau\biggl\{-\sum_{\bm{rr}^{\prime}}\psi_{\bm{r}}^{\dagger}(\tau)\bigl[\mathcal{G}^{-1}(\bm{r}-\bm{r}^{\prime},\tau)-\Sigma(\bm{r},\bm{r}^{\prime},\tau)\bigr]\psi_{\bm{r}^{\prime}}(\tau)+\sum_{\bm{q}}\frac{|\phi_{\bm{q}}(\tau)|^{2}}{2NV_{\bm{q}}}\biggr\}+\sum_{\langle\bm{rr}^{\prime}\rangle}\frac{2|\Delta_{\bm{r}-\bm{r}^{\prime}}|^{2}}{TJ},$ (8)
@@
$\Sigma(\bm{r},\bm{r}^{\prime},\tau)$ $=\Bigl\{\mathrm{i}\bigl[\partial_{\tau}\theta_{\bm{r}}(\tau)+\phi_{\bm{r}}(\tau)\bigr]\delta_{\bm{rr}^{\prime}}-t_{\bm{r}-\bm{r}^{\prime}}\bigl[\mathrm{e}^{-\mathrm{i}(\theta_{\bm{r}}(\tau)-\theta_{\bm{r}^{\prime}}(\tau))\tau_{3}}-1\bigr]\Bigr\}\tau_{3},$ (10)
@@
$+\frac{2}{NTJ}\sum_{\bm{k}\bm{q}}\bigl[\cos(k_{x}-q_{x})+\cos(k_{y}-q_{y})\bigr]\Delta_{\bm{k}}\Delta_{\bm{q}}^{*}+\sum_{\bm{q}n}\frac{|\phi_{\bm{q}n}|^{2}}{2NTV_{\bm{q}}},$ (11)
@@
$S_{\mathrm{eff}}=-\mathrm{Tr}\bigl[\ln\bigl((T\mathcal{G})^{-1}\bigr)\bigr]+\sum_{\gamma=1}^{\infty}\frac{1}{\gamma}\mathrm{Tr}\bigl[(\mathcal{G}\Sigma)^{\gamma}\bigr]+\frac{2N}{TJ}\bigl(|\Delta_{x}|^{2}+|\Delta_{y}|^{2}\bigr)+\sum_{\bm{q}n}\frac{|\phi_{\bm{q}n}|^{2}}{2NTV_{\bm{q}}},$ (12)
@@
$\Sigma(\bm{k},\bm{q},\mathrm{i}\nu_{n})\approx\frac{1}{\sqrt{N}}\bigl[\nu_{n}\tau_{3}+\mathrm{i}(\xi_{\bm{k}}-\xi_{\bm{q}})\bigr]\theta_{\bm{k}-\bm{q},n}-\frac{1}{2N}\sum_{\bm{k}^{\prime}m}(\xi_{\bm{k}}+\xi_{\bm{q}}-2\xi_{\bm{k}-\bm{k}^{\prime}})\tau_{3}\theta_{\bm{k}^{\prime}m}\theta_{\bm{k}-\bm{q}-\bm{k}^{\prime},n-m}+\frac{\mathrm{i}}{N}\phi_{\bm{k}-\bm{q},n}\tau_{3}.$ (13)
@@
$\mathrm{Tr}\bigl[(\mathcal{G}\Sigma)^{2}\bigr]\approx{}$ $\frac{1}{N}\sum_{\bm{k}\bm{q}mn}\mathrm{Tr}\Bigl[\mathcal{G}(\bm{k},\mathrm{i}\omega_{n})\bigl[\bigl(\nu_{m}\theta_{\bm{q}m}+N^{-1/2}\mathrm{i}\phi_{\bm{q}m}\bigr)\tau_{3}+\mathrm{i}(\xi_{\bm{k}}-\xi_{\bm{k}-\bm{q}})\theta_{\bm{q}m}\bigr]$
$\times\mathcal{G}(\bm{k}-\bm{q},\mathrm{i}\omega_{n-m})\bigl[\bigl(-\nu_{m}\theta_{-\bm{q},-m}+N^{-1/2}\mathrm{i}\phi_{-\bm{q},-m}\bigr)\tau_{3}-\mathrm{i}(\xi_{\bm{k}}-\xi_{\bm{k}-\bm{q}})\theta_{-\bm{q},-m}\bigr]\Bigr].$ (15)
Since $\theta_{\bm{r}}(\tau)$ is varying slowly in space and time, $\theta_{\bm{q}n}$ would quickly vanish as $\bm{q}$ and $n$ depart from zero. We can thus further expand Eqs. (14, 15) in powers of $\bm{q}$ and $\nu_{n}$ up to their quadratic order because they are controlled by $\theta_{\bm{q}n}$, that is, expand $\xi_{\bm{k}}-\xi_{\bm{k}-\bm{q}}$ to the quadratic order of $\bm{q}$ in Eq. (14) and replace $\mathcal{G}(\bm{k}-\bm{q},\mathrm{i}\omega_{n-m})$ with $\mathcal{G}(\bm{k},\mathrm{i}\omega_{n})$ and expand $\xi_{\bm{k}}-\xi_{\bm{k}-\bm{q}}$ to the linear order of $\bm{q}$ in Eq. (15). We obtain
$\mathrm{Tr}[\mathcal{G}\Sigma]$ $\approx\frac{1}{2T}\sum_{\bm{q}n}\sum_{\alpha,\beta=x,y}D^{\prime}_{\alpha\beta}q_{\alpha}q_{\beta}|\theta_{\bm{q}n}|^{2},$ (16)
$\mathrm{Tr}\bigl[(\mathcal{G}\Sigma)^{2}\bigr]$ $\approx\frac{1}{T}\sum_{\bm{q}n}\biggl[-\chi\biggl(\nu_{n}\theta_{\bm{q}n}+\frac{\mathrm{i}\phi_{\bm{q}n}}{\sqrt{N}}\biggr)\biggl(-\nu_{n}\theta_{\bm{q}n}^{*}+\frac{\mathrm{i}\phi_{\bm{q}n}^{*}}{\sqrt{N}}\biggr)+\sum_{\alpha,\beta=x,y}D^{\prime\prime}_{\alpha\beta}q_{\alpha}q_{\beta}|\theta_{\bm{q}n}|^{2}\biggr],$ (17)
@@
Here the frequency summations have been completed and $n_{F}(\cdot)$ is the Fermi distribution function and $n_{F}^{\prime}(\cdot)$ represents its derivative. We can integrate Eq. (18) by parts and use the periodicity of $\xi_{\bm{k}}$ to get the phase stiffness
@@
Finally, plugging Eqs. (16, 17) into Eq. (12) and integrating out the fields $\phi_{\bm{q}n}$ and $\theta_{\bm{q}n}$, we obtain the free energy as in Eq. (3).